OpenStudy (anonymous):

4X^2=16X-14 SOLVE THE QUADRATIC EQUATION BY COMPLETING THE SQUARE

OpenStudy (anonymous):

4x^2-16x=-14 divide by 4 x^2-4x=-14/4 Add half of the x coefficient squared to both sides x^2-4x+4=-14/4+4 (x-2)^2=-14+16/4 get the square x-2=sqrt(2/4) take square root on both sides x-2=+1/sqrt(2) x-2=-1/sqrt(2) x=2+1/sqrt(2) x=2-1/sqrt(2) done!